(3x^1.25) -4=1 , what is x?
Answer by stanbon(75887)

(3x^1.25) -4=1 , what is x?
---
3x^1.25 = 5
---
x^1.25 = 5/3
---
x^(5/4) = 5/3
---
[x^(5/4)]^(4/5) = (5/3)^(4/5)
----
x = 1.5048
================
Cheers,
Stan H.
================